OpenStudy (anonymous):

Help please!! A ball is tossed into the air from a window. Its path is described by y=-16x^2 +8x+12 where y represents the height of the ball in feet above the ground x seconds after it is tossed. What is the vertex form of the equation? What is the maximum height of the ball? *this is what I have... y=-16x^2 +8x+12 Find the x-coordinate of the vertex: x=- b/2a = 1/4 vertex form: y=-16 (x- 1/4)^2 The maximum height is 12 feet. Thanks in advance!!

OpenStudy (anonymous):

I know I am wrong I just don't know how...

OpenStudy (anonymous):

@wolf1728 do you mind helping me again?

OpenStudy (wolf1728):

Gee, I just might be able to help you.

OpenStudy (anonymous):

?

OpenStudy (wolf1728):

The vertex is at x = 1/4

OpenStudy (wolf1728):

Y value of vertex is we plug in the value of .25 y =-16x^2 +8x+12 y= -16*.25*.25 +8*.25 +12 y= -1 +2 +12 y= 13

OpenStudy (anonymous):

So the maximum height is 13 feet... right?

OpenStudy (wolf1728):

No I thought it was 12 feet (which you also calculated).

OpenStudy (anonymous):

I thought I was wrong, though..

OpenStudy (wolf1728):

Just going by the formula, y represents the height of the ball after x seconds. So, at x = 0 seconds is the height immediately after it is thrown. So, when x=0, the x terms equal zero and you are just left with 12.

OpenStudy (anonymous):

Okay, I'm kinda confused. But thank you so much! You are a big help.

OpenStudy (wolf1728):

u r welcome. Think I will have to logout for a bit - you'll be okay right?

OpenStudy (anonymous):

Yes, thank you. :)

OpenStudy (wolf1728):

Okay c ya later.